OpenStudy (anonymous):

determining the domain of a function s=sqrt3t+12

OpenStudy (anonymous):

can you show your work

OpenStudy (accessdenied):

The domain would be the values of t that we are able to use in the equation to have a real output for s. I'm unsure if you mean \[ \sqrt{3t} + 12 \] or \[ \sqrt{3t + 12} \] Either way, we'll be taking the value under the square root to be greater-than or equal-to 0.

OpenStudy (anonymous):

it is the second one you wrote.

OpenStudy (anonymous):

\[\sqrt{3t+12}\]

OpenStudy (anonymous):

I got that far and then I am not sure what to do next?

OpenStudy (accessdenied):

okay. So, we'd take "3t + 12" to be greater-than or equal to 0. (If this expression is less-than 0, then we wouldn't have a real output) 3t + 12 >= 0 3t >= -12 t >= -4 -4 <= t < infinity Interval notation: [-4, infinity)

OpenStudy (anonymous):

thank you I have one more that is on this same line but a cube root.